OpenStudy (anonymous):

Prove that if a≡b mod n and c≡d mod n, then ac≡bd mod n. a=b+kn where k is any integer c=d+ln where l is any integer (b+kn)(d+ln) n(b+k)(d+l)=n(b+d)(d+l)????

OpenStudy (anonymous):

\[ ac=(b+kn)(d+ln) = n(dk+l(b+kn)) +bd \]\[ \implies ac≡bd \pmod n \]

OpenStudy (anonymous):

Yeah dunno what you did there?

OpenStudy (experimentx):

dk+l(b+kn) is an integer ... say z ac = nz + bd

OpenStudy (anonymous):

stlil confused.

OpenStudy (experimentx):

a = b + nk ... for some integer k, it implies a≡b mod n for some integer z ac = nz + bd ac≡bd mod n

OpenStudy (anonymous):

@experimentX how is dk+l(b+kn) an integer??

OpenStudy (experimentx):

because every element is an integer ,,